OpenStudy (anonymous):

solve 3^(x-1)= (1/9)

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

how though?

OpenStudy (anonymous):

3^(x-1)=9^-1 3^(x-1)=3^-2 compare powers x-1=-2 x=-1

OpenStudy (anonymous):

suppose equation =y....then separate x value..

OpenStudy (anonymous):